Find both the -intercept and the -intercept of the line given by the equation
8.2x +7.2y-7.7=0

Round your answers to decimal places

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Find both the -intercept and the -intercept of the line given by the equation
8.2x +7.2y-7.7=0
8.2x + 7.2y = 7.7
-------
I suppose you meant x and y intercepts.
For x intercept, set y = 0 and vice versa.
x = 0: 7.2y = 7.7
y = 7.7/7.2 = 1.06944444...
---------
y = 0: 8.2x = 7.7
x = 0.9390244

Answer by stanbon(75887)

Find both the -intercept and the -intercept of the line given by the equation
8.2x +7.2y-7.7=0
----------------------------
To find the y-intercept.
Let x = 0
7.2y = 7.7
y = 1.069444..
----------------
To find the x-intercept.
Let y = 0
8.2x = 7.7
x = 0.93902..
==================
Cheers,
Stan H.